# Training-Free Diffusion Priors for Text-to-Image Generation via Optimization-based Visual Inversion

### v1

@@ 6 Conclusions and Future Work
In this report, we presented an initial exploration of Optimization-based Visual Inversion (OVI) as a training-free diffusion prior in the context of text-to-image generation. Our experiments suggest that this approach could be a viable alternative to traditional expensive trained priors and, in its unconstrained form, helped expose a critical weakness in current compositional T2I benchmarks. Directly feeding the text embedding into the decoder often exceeds sota expensive priors in the evaluated benchmark. Additionally we show that by incorporating Mahalanobis and Nearest-Neighbor constraints, we can steer the OVI-based optimization process to produce higher visual fidelity images. Our Nearest-Neighbor constrained OVI, in particular, achieves a better quantitative score and competitive visual quality compared to the state-of-the-art ECLIPSE prior, all without requiring any prior training.
For future work, we see several promising directions. The Nearest-Neighbor constraint could be enhanced by using a K-Nearest Neighbors approach, averaging the embeddings of several close neighbors to provide a more robust target. A Gaussian Mixture Model could also be used to model the distribution of image embeddings more accurately than a single Gaussian. Most importantly, our work underscores the need for better evaluation metrics that capture the full spectrum of image quality, including photorealism, coherence, and compositional correctness, beyond simple semantic similarity.

### v3

@@ 6 Conclusions and Future Work
In this work, we investigated Optimization-based Visual Inversion (OVI) as a training-free diffusion prior in the context of text-to-image generation. Our experiments indicate that this approach is a viable alternative to traditional expensive trained priors and, in its unconstrained form, helped expose a critical weakness in current compositional T2I benchmarks. Directly feeding the text embedding into the decoder often outperforms expensive state-of-the-art priors in the evaluated benchmark. Additionally, we show that by incorporating Mahalanobis and Nearest-Neighbor constraints, we can steer the OVI-based optimization process to produce higher visual fidelity images. Our Nearest-Neighbor constrained OVI, in particular, achieves a better quantitative score and competitive visual quality compared to the leading ECLIPSE prior, all without requiring any prior training and utilizing a reference set orders of magnitude smaller ($\approx$0.8% of the latter’s training data).
For future work, we see several promising directions. The Nearest-Neighbor constraint could be enhanced by using a K-Nearest Neighbors approach, averaging the embeddings of several close neighbors to provide a more robust target. A Gaussian Mixture Model bishop2006pattern could also be used to model the distribution of image embeddings more accurately than a single Gaussian. Most importantly, our work underscores the need for better evaluation metrics that capture the full spectrum of image quality, including photorealism, coherence, and compositional correctness, beyond simple semantic similarity.